OpenStudy (anonymous):

Use the Pythagorean Theorem to determine if 1, √3, and 2 could be the sides of a right triangle. Explain your answer.

OpenStudy (ranga):

If the three sides satisfy the condition, a^2 = b^2 + c^2, then it is a right triangle. Choose the largest side as a. In this case 2 is the largest of 1, sqrt(3) and 2. a= 2, b = 1, c = sqrt(3) Does it satisfy a^2 = b^2 + c^2?